# The Fragile Nature of Road Transportation Systems

### v2

@@ Abstract
Major cities worldwide experience problems with the performance of their road transportation systems, and the continuous increase in traffic demand presents a substantial challenge to the optimal operation of urban road networks and the efficiency of traffic control strategies. Although robust and resilient transportation systems have been extensively researched over the past decades, their performance under an ever-growing traffic demand can still be questionable. The operation of transportation systems is widely believed to display fragile property, i.e., the loss in performance increases exponentially with the linearly increasing magnitude of disruptions, which undermines their continuous operation. Nowadays, the risk engineering community is embracing the novel concept of antifragility, which enables systems to learn from historical disruptions and exhibit improved performance as disruption levels reach unprecedented magnitudes. In this study, we demonstrate the fragile nature of road transportation systems when faced with demand or supply disruptions. First, we conducted a rigorous mathematical analysis to establish the fragile nature of the systems theoretically. Subsequently, by taking into account real-world stochasticity, we implemented a numerical simulation with realistic network data to bridge the gap between the theoretical proof and the real-world operations, to reflect the potential impact of uncertainty on the fragile property of the systems. This work aims to demonstrate the fragility of road transportation systems and help researchers better comprehend the necessity to explicitly consider antifragile design for future traffic control strategies, coping with the constantly growing traffic demand and subsequent traffic accidents.
@@
(anti-)fragility, road transportation systems, macroscopic fundamental diagram, model stochasticity
@@ 1 Introduction
As reported by both the U.S. Department of Transportation (2019) and Federal Statistical Office of Switzerland (2020), motorized road traffic before the pandemic has experienced an approximate 50% growth over the past few decades. Researchers have also found that the continuous growth in traffic volume has consequently contributed to a rise in disruptive events, such as severe congestion and more frequent accidents (Dickerson et al., 2000; Chang and Xiang, 2003). With evidence confirming the recovery of individual motorized traffic from the pandemic (Büchel et al., 2022; Marra et al., 2022; Ciuffini et al., 2023), it is expected that this upward trend will continue in the coming decades (Zhang and Zhang, 2021).
@@
Meanwhile, there is a common understanding that road transportation networks can exhibit fragile properties. Fragility signifies a system’s susceptibility to exponentially escalating performance deterioration as disruptions increase in magnitude. One prominent example of such fragile characteristics is the BPR function (U.S. Bureau of Public Roads, 1964), which distinctly illustrates with empirical data at the link level that travel time grows exponentially with traffic flow, leading to an infinite temporal cost when the traffic influx is at the maximal density of the network. Moreover, not limited to urban road networks, many other types of transportation systems can also display similar fragile responses to increasing disruption levels, such as in railway (Gibson et al., 2002; Saidi et al., 2023) and aviation systems (U.S. Congress, Office of Technology Assessment, 1984).
@@
Researchers have devoted extensive efforts to the assessment and design of robust and resilient transportation systems of all kinds, such as in railway systems (Corman et al., 2014; Larsen et al., 2014), public transportation operation (Cats, 2016; Fuchs and Corman, 2019), aviation (Isaacson et al., ), and road networks (Ampountolas et al., 2017; Yang et al., 2019; Leclercq et al., 2021). However, when accounting for the ever-growing traffic volume in urban road networks and the exponentially escalating adversarial consequences, it is natural for us to wonder whether the current level of robustness and resilience can still guarantee the performance of road networks considering a long-term time horizon. Hence, we introduce the cutting-edge concept of antifragility to explain the phenomenon that roadside performance deteriorates exponentially with linearly growing disruptions. Previous studies discussing the fragile response, such as the BPR function, have primarily relied on empirical data and intuitive reasoning rather than rigorous mathematical proof. This paper, on the other hand, serves as a proof of concept, aiming to establish the fragile nature of road transportation systems through mathematical analysis. Additionally, as stochasticity prevails in transportation systems in the real world, we also designed a numerical simulation considering real-world stochasticity, to study to what extent such realistic uncertainties can influence the fragile characteristics of transportation systems. The overarching objective of this paper is to provide insights to transportation researchers for the future design of transportation systems and control strategies to be not only robust and resilient but also antifragile.
@@
The remainder of this paper is structured as follows. Section 2 introduces antifragility and reviews specific traffic-related mathematical models. Section 3 formulates the mathematical definition of (anti-)fragility and its applications. Then we conduct the mathematical proof in Section 4, whereas Section 5 presents the numerical simulation with real-world network and stochasticity. With Section 6, we conclude the fragile nature of road transportation systems and its implications for future studies.
@@ 2 Relevant Literature
Various terminologies have been proposed to evaluate the performance of road transportation systems, including their control strategies, and two commonly used terms to characterize the extent of performance variations under stress are robustness (Duan and Lu, 2014; Shang et al., 2022) and resilience (Mattsson and Jenelius, 2015; Calvert and Snelder, 2018). However, the definitions of robustness and resilience can vary under different contexts, even within the transportation domain itself, and are sometimes used interchangeably due to the different backgrounds and points of view of the authors (Tang et al., 2020; Corman et al., 2018). In this study, we adopt the definition proposed in Zhou et al. (2019), wherein robustness involves evaluating a system’s ability to maintain its initial state and withstand performance degradation when confronted with uncertainties and disturbances. On the other side, resilience emphasizes a system’s capability and promptness in recovering from major disruptions and returning to its original state. In brief, robustness relates to resistance, whereas resilience is about recovery.
@@
Nevertheless, these two terms can overlook the consideration of a longer timespan and the potential escalation of disruptions, which is particularly relevant in transportation as the traffic demand grows continuously and consequently so as the occurrences of accidents. Thus, there is the necessity to introduce a new term to address this gap. The novel concept of antifragility was initially proposed in Taleb (2012) and mathematically elaborated in Taleb and Douady (2013); Taleb and West (2023), and it serves as a general concept aimed at transforming people’s understanding and perception of risk. By embracing current risks, we can potentially leverage and adapt to future risks of greater magnitudes. When employed in systems and control, (anti-)fragility can be conceptualized as a nonlinear relationship between the performance and the magnitude of disruptions. If the performance is compromised due to unexpected disruptions, the relationship between the loss in performance and the disruptions would be convex for a fragile system, while being concave for an antifragile system. Ever since being proposed, antifragility has gained popularity in the risk engineering community across multiple disciplines, such as economy (Manso et al., 2020), biology (Kim et al., 2020), medicine (Axenie et al., 2022), energy (Coppitters and Contino, 2023), robotics (Axenie and Saveriano, 2023), and lately in transportation (Sun et al., 2024). It should also be highlighted that although systems can be fragile by nature, proper intervention and control strategies can enhance their antifragility against increasing disruptions (Axenie et al., 2023).
@@
In transportation, by unveiling the convex relationship between travel time and traffic flow with empirical data, the BPR function (U.S. Bureau of Public Roads, 1964), as shown in Eq. 1 has given an intuitive example showing the fragility of the transportation systems on a link level. Together with its variations (Dowling and Skabardonis, 1993; Skabardonis and Dowling, 1997), the BPR function has been extensively applied in the estimation of the link (route) travel time (Lo et al., 2006; Ng and Waller, 2010; Wang et al., 2014). However, to uphold the statement that road transportation systems are fragile in general, an empirical function like BPR alone is not sufficient without rigid mathematical proof. It is also desired to show the fragility more broadly, i.e., not only on the link level but also on the macroscopic level, as well as for different types of disruptions.
$T$ $=T_{ff}\left(1+\alpha\left(\frac{q}{q_{\rm max}}\right)^{\beta}\right)$ (1)
@@
The assessment of traffic performance can be conducted at either the microscopic or macroscopic level, sometimes also mesoscopically (Zehe et al., 2015). However, due to the non-identical traffic characteristics across different levels, researchers have formulated diverse models to offer a more precise description of traffic dynamics on different levels. While some models are generated numerically from data, others are derived through analytical methods (Mariotte et al., 2017).
@@
The earliest study on traffic performance was carried out and described in Greenshields et al. (1934) on a section of a highway and yielded the first Fundamental Diagram (FD), as shown in Fig. 1(a), which exhibits the relationship between traffic flow, denoted as $q$, and density, denoted as $k$, in the shape of a second-degree polynomial. Later, other researchers also developed FDs in various forms, with one of the most commonly applied FDs being proposed in Daganzo (1994), as shown in Fig. 1(b), which is characterized by two linear functions. While the Greenshields FD is entirely generated through fitting the polynomial coefficients with on-site data, the Daganzo FD, on the other hand, is derived analytically by treating the traffic flow hydrodynamically and incorporates variables with physical meanings, i.e., the free-flow speed, back-propagation speed, and critical density, denoted as $v_{f}$, $w$, and $k_{c}$, respectively.
@@
On the network level, with the assumption of a homogeneous region, a Macroscopic Fundamental Diagram (MFD) can be produced by aggregating data points gathered from representative links within this network. Similar to FDs, MFDs can be approximated through different models, and the most widely applied approaches include polynomials and multi-regime linear functions. MFDs can be fit numerically from field measurements as a cubic polynomial, as in Haddad and Shraiber (2014); Sirmatel and Geroliminis (2018) and illustrated in Fig. 2(a), and therefore, it’s commonly seen in research works simulating traffic control on a macroscopic level with a given MFD. On the other hand, based on the framework of variation theory (Daganzo, 2005), Daganzo and Geroliminis (2008) is the first study to analytically generate an MFD, and is often referred to as the Method of Cuts (MoC). Instead of installing loop detectors under the pavement and gathering massive traffic flow data, MoC can be applied to derive an MFD directly from roadside variables with physical meanings, such as free flow speed, traffic signal cycle, lane length, etc. An example of multi-regime linear functions MFD from MoC is shown in Fig. 2(b). Leclercq and Geroliminis (2013) modified the original MoC to accommodate topology and signal timing heterogeneity within the network. Although theoretically an infinite number of cuts can be generated to approximate the MFD, only the practical cuts, the solid lines as shown in Fig. 2(b), are used for simplicity reasons. Although some other analytical methods have also been proposed to produce an MFD, such as through stochastic approximation (Laval and Castrillón, 2015), Tilg et al. (2020) has demonstrated that MoC yields a more accurate upper bound for the MFD. Lately, researchers have also explored the potential utilization of MFD in railway (Corman et al., 2019) and aviation operations (Safadi et al., 2023), which yields some positive possibilities of extending the MFD into other modes.
@@
As also acknowledged in these works, some variables, particularly the parameters related to signalization, can be hard to acquire in the real world with actuated signals. Therefore, simplified MFDs are also widely adopted when solving more applicational problems in the real world. Daganzo et al. (2018) approximated a simplified MoC-based MFD using only three cuts from one of each stationary, forward, and backward observer, and formed the uMFD. Another shortcoming of the numerical-based MFD is that, since the data points near the maximal density are scarce overall, the approximated MFD can lose its fidelity near the maximal density. For example, the third-degree polynomial in Geroliminis et al. (2013) has only one real root, indicating the speed is above zero despite the fact that the network has already reached a gridlock, which cannot be realistic. As claimed in Daganzo and Geroliminis (2008), since MFDs should be concave, the flow and speed are both zero by definition at maximal density.
@@ 3 Problem formulation
As explained in Section 2, an (anti-)fragile response of a system can be characterized through a nonlinear relationship between the performance loss and the magnitude of the disruption, as shown in Fig. 3(a). Both nonlinear functions can be represented by Jensen’s inequality (Jensen, 1906), with either $E[g(X)]\geq g(E[X])$ for a fragile response or $E[g(X)]\leq g(E[X])$ for an antifragile response. This relationship can then be determined through the second derivative (Ruel et al., 1999), i.e., a positive second derivative featuring a convex function and hence a fragile system and vice versa. It should be noted that the calculation of the derivatives is only possible when the function is continuous and differentiable, which means the underlying mathematical model representing the system needs to be known beforehand.
However, in most real-world scenarios, the mathematical function of the system is agnostic, and only discrete measurements of the system’s performance are available. In this case, we can calculate the distribution skewness to determine the (anti-)fragile property of the system, as illustrated with an example of financial deficit in Taleb and Douady (2013) and Coppitters and Contino (2023) when designing an antifragile renewable energy system. A negative skewness represents the long tail pointing to the left and indicates an antifragile response, as shown in Fig. 3(b).
(a) Nonlinear relationship for a continuous function
(b) Distribution skewness for discrete measurements
Figure 3: Characteristics and identification of (anti-)fragility
In this paper, we address three sets of opposing concepts for the mathematical analysis when studying the fragile nature of road transportation systems:
    microscopic / macroscopic
@@
As microscopic and macroscopic traffic models were introduced in the previous Section 2, here we start with differentiating between demand and supply disruptions, with the latter sometimes referred to as MFD disruptions. Since traffic networks primarily involve the management of supply and demand, we consider that any traffic disruption in the real world can be classified as either a demand or a supply disruption. A demand disruption can be easily understood as, for example, surging traffic due to a social event, whereas a supply disruption may indicate an impaired network due to external factors, such as adversarial weather or lane closure. Additionally, since disruptions represent abnormal cases that only exist temporarily, we consider not only the onset of disruptions but also assess the recovery process of the systems following such disruptions. This coincides with the aforementioned definitions of robustness and resilience, i.e., robustness is about resistance against disruptions while resilience is about the recovery from them. By considering both the onset and recovery processes, it becomes possible to compare the performance between either robust and antifragile designs or resilient and antifragile designs. The scheme of onset and recovery from demand or supply disruptions are illustrated in Fig. 4(a) and Fig. 4(b), and since FDs share similar profiles as MFDs, we simply use MFDs here as examples for illustration. We denote the FD/MFD profile as $G(k)$ and assume a constant base demand in the network as $q$, resulting in an equilibrium traffic state in the network in the absence of any disruption. The initial density at equilibrium, the critical density, the new density after disruption, and the gridlock density are denoted as $k_{0}$, $k_{c}$, $k^{\prime}$, and $k_{\rm max}$, respectively. For the study of supply disruptions, we introduce a disruption magnitude coefficient, denoted as $r$, so that the disrupted MFD profile can be represented as $(1-r)G(k)$. It should also be noted that on the network level, instead of traffic flow - density, MFD can also be represented with vehicle accumulation - trip completion, such as in Zhou and Gayah (2021); Kouvelas et al. (2017); Genser and Kouvelas (2022). Hence, with vehicle accumulation denoted as $n$, $k_{*}$ and $G(k)$ can also be replace with $n_{*}$ and trip completion $G(n)$.
@@
Several assumptions need to be established to define the scope of our work. Also, as the study of fragility involves the recovery process from disruptions, a critical condition to avoid here is the network succumbing to a complete gridlock, where recovery is not possible anymore.

### v4

@@ Abstract
Major cities worldwide experience problems with the performance of their road transportation systems, and the continuous increase in traffic demand presents a substantial challenge to the optimal operation of urban road networks and the efficiency of traffic control strategies. The operation of transportation systems is widely considered to display fragile property, i.e., the loss in performance increases exponentially with the linearly increasing magnitude of disruptions. Meanwhile, the risk engineering community is embracing the novel concept of antifragility, enabling systems to learn from historical disruptions and exhibit improved performance under black swan events. In this study, based on established traffic models, namely fundamental diagrams and macroscopic fundamental diagrams, we first conducted a rigorous mathematical analysis to prove the fragile nature of road transportation systems theoretically. Subsequently, we propose a skewness-based indicator that can be readily applied to cross-compare the degree of fragility for different networks solely dependent on the MFD-related parameters. At last, by taking real-world stochasticity into account, we implemented a numerical simulation with realistic network data to bridge the gap between the theoretical proof and the real-world operations, to reflect the potential impact of uncertainty on the fragility of the systems. This work aims to demonstrate the fragile nature of road transportation systems and help researchers better comprehend the necessity to consider explicitly antifragile design for future traffic control strategies.
@@
antifragility , traffic disruptions , road transportation systems , macroscopic fundamental diagram , fragility indicator
journal: Elsevier
@@ 1 Introduction
As reported by both the U.S. Department of Transportation (2019) and the Federal Statistical Office of Switzerland (2020), motorized road traffic before the pandemic has experienced an approximate 50% growth over the past few decades. Researchers have also found that the continuous growth in traffic volume has consequently contributed to a rise in disruptive events, such as severe congestion and more frequent accidents (Dickerson et al., 2000). With evidence confirming the recovery of individual motorized traffic from the pandemic (Büchel et al., 2022; Marra et al., 2022), it is expected that this upward trend will continue in the coming decades, even with the political and behavioral shift being taken into account, such as bicycle-friendly design and working-from-home (Zhang and Zhang, 2021). Matthias et al. (2020) shows that although Germany is widely regarded as an active player in carbon neutrality, motorized traffic is still projected to increase following a more politically regulated sustainable development. Another emerging transformation often disregarded is the deployment of Autonomous Vehicles (AVs) and autonomous mobility-on-demand, which may account for more than 10% growth in terms of induced demand(Nahmias-Biran et al., 2021). Therefore, future road networks are highly expected to experience a further increase in traffic demand.
@@
Meanwhile, there is a common understanding that road transportation networks can exhibit fragile properties. Fragility signifies a system’s susceptibility to exponentially escalating performance deterioration as disruptions increase in magnitude. One prominent and intuitive example of such fragile characteristics is the BPR function (U.S. Bureau of Public Roads, 1964), which distinctly illustrates with empirical data at the link level that travel time grows exponentially with traffic flow, leading to an infinite temporal cost when the traffic influx is at the maximal density of the network. Moreover, not limited to urban road networks, many other types of transportation systems can also display similar fragile responses to increasing disruption levels, such as in railway (Saidi et al., 2023) and aviation systems (U.S. Congress, Office of Technology Assessment, 1984).
@@
Researchers have devoted extensive efforts to the assessment and design of robust and resilient transportation systems of all kinds, such as in railway systems (Corman et al., 2014), public transportation operation (Cats, 2016), aviation (Isaacson et al., 2010), and road networks (Ampountolas et al., 2017; Yang et al., 2019; Leclercq et al., 2021). However, when accounting for the ever-growing traffic volume in urban road networks and the exponentially escalating adversarial consequences, it is natural for us to wonder whether the current level of robustness and resilience can still guarantee the performance of road networks considering a long-term time horizon. Hence, we introduce the cutting-edge concept of antifragility to explain the phenomenon that urban road performance deteriorates exponentially with linearly growing disruptions. Previous studies on traffic performance showing signs of fragility, such as the BPR function, have primarily relied on empirical data and intuitive reasoning rather than rigorous mathematical proof. This paper, on the other hand, makes the following contributions:
  1. 1.
    Proof of concept: This research aims to establish the fragile nature of road transportation systems with a framework through mathematical analysis considering different aspects, that are at the link and the network levels, under demand and supply disruption, as well as considering the onset of disruption and recovery from disruption.
  2. 2.
    Methodological contribution: A skewness-based fragility indicator inspired by the Sigmoid curve is developed for the approximation of the fragility of a specific network. A scalable unit MFD can be applied for the cross-comparison of the fragility among different networks, relying on merely the parameters with physical meanings.
  3. 3.
    From theory to reality: As stochasticity prevails in road transportation systems in the real world, we also designed a numerical simulation considering real-world stochasticity, to study to what extent such realistic uncertainties can influence the fragile characteristics of transportation systems.
The overarching objective of this paper is to provide insights to transportation researchers for the future design of transportation systems and control strategies to be not only robust and resilient but also antifragile.
@@
The remainder of this paper is structured as follows. Section 2 introduces antifragility and reviews specific traffic-related mathematical models. Section 3 formulates the mathematical definition of (anti-)fragility and its detection methods. Then we conduct the mathematical proof to establish such fragile nature in Section 4, whereas a skewness-based fragility indicator is proposed in Section 5. Section 6 presents the numerical simulation with real-world network and stochasticity. With Section 7, we conclude the fragile nature of road transportation systems and its implications for future studies.
@@ 2 Relevant Literature
Various terminologies have been proposed to evaluate the performance of road transportation systems, including their control strategies, and two commonly used terms to characterize the extent of performance variations under stress are robustness (Shang et al., 2022) and resilience (Mattsson and Jenelius, 2015; Calvert and Snelder, 2018). However, the definitions of robustness and resilience can vary under different contexts, even within the transportation domain itself, and are sometimes used interchangeably due to the different backgrounds and points of view of the authors (Corman et al., 2018; Tang et al., 2020). In this study, we adopt the definition proposed in Zhou et al. (2019), wherein robustness involves evaluating a system’s ability to maintain its initial state and withstand performance degradation when confronted with uncertainties and disturbances. On the other side, resilience emphasizes a system’s capability and promptness to recover from major disruptions and return to its original state. In brief, robustness relates to resistance, whereas resilience is about recovery. Another very recent term attracting much attention is the self-organized criticality of traffic flow from a complex systems perspective (Laval, 2023), implying an avalanche effect of unforeseeable magnitude when traffic ever becomes critical, which resembles the philosophy of optimality versus fragility.
@@
Nevertheless, both robustness and resilience can overlook the consideration of a longer timespan, which is particularly relevant in transportation as the traffic demand grows continuously and consequently so as the occurrences of accidents. Thus, it is necessary to introduce a new term to address this gap. The novel concept of antifragility was initially proposed in Taleb (2012) and mathematically elaborated in Taleb and Douady (2013), and it serves as a general concept aimed at transforming people’s understanding and perception of risk. By embracing current risks, we can potentially leverage and adapt to future risks of greater magnitudes. When employed in systems and control, (anti-)fragility can be conceptualized as a nonlinear relationship between the performance and the magnitude of disruptions. If the performance is compromised due to unexpected disruptions, the relationship between the loss in performance and the disruptions would be convex for a fragile system, while being concave for an antifragile system. Ever since being proposed, antifragility has gained popularity in the risk engineering community and across multiple disciplines, such as biology (Kim et al., 2020), medicine (Axenie et al., 2022), energy (Coppitters and Contino, 2023), robotics (Axenie and Saveriano, 2023), and lately in transportation (Sun et al., 2024). It should also be highlighted that although systems can be fragile by nature, proper intervention and control strategies can enhance their antifragility against increasing levels of disruptions (Axenie et al., 2024).
@@
For each link in a highway network, the BPR function (U.S. Bureau of Public Roads, 1964), as shown in Eq. 1, has given an intuitive example showing the fragility of the transportation systems at a link level. As the traffic flow $q$ approaches the capacity of a link $q_{\rm max}$, the travel time $T$ is revealed to grow exponentially compared to the free flow travel time $T_{ff}$. The BPR function and its variations have been extensively applied in the estimation of the link (route) travel time (Lo et al., 2006; Ng and Waller, 2010; Wang et al., 2014). However, to uphold the statement that road transportation systems are fragile in general, an empirical function like BPR alone is not sufficient without rigid mathematical proof. It is also desired to show the fragility more broadly, i.e., not only at the link level but also at the network level, as well as for different types of disruptions.
$T$ $=T_{ff}\left(1+0.15\left(\frac{q}{q_{\rm max}}\right)^{4}\right)$ (1)
@@
The assessment of the traffic performance of road transportation systems can be conducted at either the link or the network level. However, due to the non-identical traffic characteristics across different levels, researchers have formulated diverse models to offer a more precise description of traffic dynamics at these levels, demonstrating the mathematical relationship between traffic flow, denoted as $q$, traffic density, denoted as $k$, and sometimes also traffic speed, denoted as $v=\frac{q}{k}$. While some traffic models are generated numerically from data collected with sensors such as loop detectors or aerial photos, others are derived through analytical methods (Mariotte et al., 2017). Typically, these models also involve a critical density, denoted as $k_{c}$, below which the flow increases as density rises, but beyond which the flow decreases as density continues to increase until reaching the maximal density, denoted as $k_{\rm max}$.
@@
The earliest study on traffic performance was carried out and described in Greenshields et al. (1934) at a section of a highway and yielded the first Fundamental Diagram (FD), as shown in Fig. 1(a) exhibiting the relationship between traffic flow and density at the link level in the shape of a second-degree polynomial. Later, other researchers also developed FDs in various forms, with one of the most commonly applied FDs being proposed in Daganzo (1994), as shown in Fig. 1(b), which is characterized by two linear functions. While the Greenshields FD is entirely generated by fitting the polynomial coefficients with on-site data, the Daganzo FD, on the other hand, is derived analytically by treating the traffic flow hydrodynamically and incorporates variables with physical meanings, namely the gradient of the line segment below $k_{c}$ as the free-flow speed, represented by $u_{f}$, and the gradient of the line segment beyond $k_{c}$ as the backward wave speed, represented by $w$, a virtual wave of reduced speed and increased density propagating against the direction of traffic flow when vehicles are forced to stop.
@@
On the network level, with the assumption of a homogeneous region, a Macroscopic Fundamental Diagram (MFD) can be produced by aggregating data points gathered from representative links within a network. Similar to FDs, MFDs can be approximated through different models, and the most widely applied functional forms include polynomials and multi-regime linear functions. MFDs can be fit numerically from field measurements as a cubic polynomial, as in Haddad and Shraiber (2014); Sirmatel and Geroliminis (2018) and illustrated in Fig. 2(a). Therefore, it’s commonly seen in research works simulating traffic control at a network level with a given MFD. On the other hand, based on variation theory (Daganzo, 2005) and the assumption to simply a homogenous network into an abstract corridor, Daganzo and Geroliminis (2008) is the first study to generate an MFD analytically, and is often referred to as the Method of Cuts (MoC). Instead of installing loop detectors under the pavement and gathering massive traffic flow data, MoC can be applied to derive an MFD directly from traffic-related variables with physical meanings, such as free flow speed, traffic signal cycle, lane length, etc. An example of multi-regime linear functions MFD from MoC is shown in Fig. 2(b). Although theoretically an infinite number of cuts can be generated to approximate the MFD, only the practical cuts, the solid lines as shown in Fig. 2(b), are used for simplicity. Leclercq and Geroliminis (2013) improved the original MoC to accommodate topology and signal timing heterogeneity within the network. Ambühl et al. (2019); Saedi et al. (2020) also introduced algorithms to partition an entire heterogeneous network into multiple more homogeneous regions. The most recent advance in the study of MFD built on MoC is Tilg et al. (2023), which relaxes the assumption of demand homogeneity and an abstract corridor by creating a hypernetwork from a set of corridors and incorporating turn ratios between corridors. Although some other analytical methods have also been proposed to produce an MFD, such as through stochastic approximation (Laval and Castrillón, 2015), Tilg et al. (2020) has demonstrated that MoC yields a more accurate upper bound for the MFD. Lately, researchers have also explored the potential utilization of MFD in railway (Corman et al., 2019) and aviation operations (Safadi et al., 2023), which yields some positive possibilities of extending the MFD into other modes.
@@
As also acknowledged in these works, some variables, particularly the parameters related to signalization, can be hard to acquire in the real world with actuated signals. Therefore, simplified MFDs are also widely adopted when solving more applicational problems in the real world. Daganzo et al. (2018) approximated a simplified MoC-based MFD using only three cuts from one of each stationary, forward, and backward observer, and formed the trapezoidal uMFD. Moreover, one shortcoming of the numerical-based MFD is that, since the data points near the maximal density $k_{\rm max}$ are scarce overall, the approximated MFD can lose its fidelity near $k_{\rm max}$. For example, the third-degree polynomial in Geroliminis et al. (2013) has only one real root, indicating the speed is above zero despite the network already reaching a gridlock, which is unrealistic. As claimed in Daganzo and Geroliminis (2008), since MFDs should be concave, the flow and speed are both zero by definition at maximal density.
The above discussion in the literature is confined to the flow-density MFD. However, assuming the total lane length and average trip length, denoted by $D$ and $L$ respectively, are known for a given network, the trip completion (sometimes also referred to as network production or outflow rate in different research works) denoted by $m$, and vehicle accumulation within the network denoted by $n$, can be further determined (Leclercq and Geroliminis, 2013; Geroliminis et al., 2013), as shown in Eq. 2a and Eq. 2b. The advantage of using an $m-n$ MFD is that it can serve as a transfer function to deduce the future system dynamics, which is particularly common practice in the study of various traffic control strategies, as in (Rodrigues and Azevedo, 2019; Zhou and Gayah, 2021; Genser and Kouvelas, 2022).
$n$ $=kD$ (2a)
$m$ $=\frac{qD}{L}$ (2b)
@@ 3 Problem formulation
In this study, we focus on the relationship between vehicle time spent versus disruption magnitudes, following a loss-disruption relationship, however, a bigger picture of (anti-)fragility as shown in Fig. 3 can help facilitate the understanding of the gist behind it. In this four-quadrant diagram, while the first quadrant shows a gain-opportunity relationship, the third quadrant illustrates the loss-disruption relationship, which is the focus of the study. Although convexity signifies an antifragile response on the whole four-quadrant diagram, when it comes to each quadrant, convexity or concavity itself does not necessarily indicate (anti-)fragility without the context. Also, a variable can be an opportunity for one indicator but becomes a disruption for another. For example, while traffic density always imposes a non-positive effect on travel speed and thus follows the loss-disruption relationship with the fragile response being convex, vehicle accumulation can actually be beneficial to trip completion so that the related fragility curve is concave. In our case, as highlighted in the beige color, both nonlinear functions can be represented by Jensen’s inequality, with either $E[g(X)]\geq g(E[X])$ for a fragile response or $E[g(X)]\leq g(E[X])$ for an antifragile response. This relationship can then be determined through the second derivative (Ruel et al., 1999), i.e., a positive second derivative featuring a convex function and hence a fragile system and vice versa. It should be noted that the calculation of the derivatives is only possible when the function is continuous and differentiable, which means the underlying mathematical model representing the system needs to be known beforehand.
Figure 3: Antifragility in a four-quadrant diagram
However, in most real-world scenarios, the mathematical function of the system is agnostic, and only discrete measurements of the system’s performance are available. In this case, we can calculate the distribution skewness to determine the (anti-)fragile property of the system, as illustrated with an example of financial deficit in Taleb and Douady (2013) and Coppitters and Contino (2023) when designing an antifragile renewable energy system. For a loss-disruption relationship, a positive skewness represented by the long tail pointing to the left indicates a fragile response, whereas a negative skewness showcases the fragility of a gain-opportunity relationship.
For the following mathematical analysis of the fragile nature of road transportation systems under a unified framework, we address three sets of opposing concepts:
    link level / network level
@@
As link-level and network-level traffic models were introduced in the previous Section 2, we start with differentiating between demand and supply disruptions. Since traffic networks primarily involve the management of supply and demand, any traffic disruption in the real world can be classified as either a demand or a supply disruption. A demand disruption can be easily understood as, for example, surging traffic due to a social event, whereas a supply disruption may indicate an impaired network due to external factors, such as adversarial weather or lane closure. Additionally, since disruptions represent abnormal cases that only exist temporarily, we consider not only the onset of disruptions but also assess the recovery process of the systems following such disruptions. This coincides with the definitions of robustness and resilience, i.e., robustness is about resistance against disruptions while resilience is about the recovery from them. By considering both the onset and the recovery processes, it is possible to compare the performance between either robust and antifragile designs or resilient and antifragile designs. The scheme of onset and recovery from demand or supply disruptions are illustrated in Fig. 4(a) and Fig. 4(b), and since FDs share similar profiles as MFDs, we use MFDs here as generic examples for illustration.
@@
We denote the FD/MFD profile as $G(k)$ and assume a constant base demand in the network as $q_{0}$, resulting in an equilibrium traffic state without any disruption. The initial density at equilibrium, the critical density, the new density after disruption, and the gridlock density are denoted as $k_{0}$, $k_{c}$, $k^{\prime}$, and $k_{\rm max}$, respectively. For the study of supply disruptions, we introduce a disruption magnitude coefficient, denoted as $r$, so that the disrupted MFD profile can be represented by $(1-r)G(k)$. It should also be noted that at the network level, MFD can also be represented with the vehicle accumulation -- trip completion relationship instead of the flow -- density relationship. Hence, with vehicle accumulation denoted as $n$, $k^{\prime}$ and $G(k)$ can also be replaced with $n^{\prime}$ and trip completion $M(n)$. For clarity, a notation list is summarized in Table 1. Another separate table of notations with real-world values solely for the numerical simulation can be found in Section 6.
Table 1: List of notations
FD/MFD-related notations
$N$ Total number of cuts according to MoC
$y$ Starting cut (the most congested cuts)
$z$ Ending cut (the least congested cuts)
$i$ Sequential number of cuts from the most to the least congested cuts
Flow - density Trip completion - accumulation
$G(\cdot)$ Functional form of FD/MFD $M(\cdot)$ Functional form of MFD
$k$ Density $n$ Vehicle accumulation
$k_{c}$ Critical density $n_{c}$ Critical vehicle accumulation
$k_{\rm max}$ Maximal density at gridlock $n_{\rm max}$ Maximal accumulation at gridlock
$k_{0}$ Initial density at equillibrium $n_{0}$ Initial accumulation at equillibrium
$k^{\prime}$ Disruption density $n^{\prime}$ Disruption vehicle accumulation
$q$ Traffic flow $m$ Trip completion
$q_{\rm max}$ Capacity of the network, $k=k_{c}$ $m_{\rm max}$ Maximal trip completion, $n=n_{c}$
$q_{0}$ Base demand flow, $k=k_{0}$ $m_{0}$ Base demand inflow, $n=n_{0}$
$q^{\prime}$ Flow on disruption density, $k=k^{\prime}$ $m^{\prime}$ Completion at disruption, $n=n^{\prime}$
$v$ Speed, $v=q/k$ $a$ Completion rate, $a=m/n$
$u_{i}$ Gradient of cut $i$ $a_{i}$ Gradient of cut $i$
$c_{i}$ Intercept of cut $i$ $b_{i}$ Intercept of cut $i$
$u_{f}$ Free flow speed, $u_{f}=u_{N}$ $a_{f}$ Gradient of the free flow cut
$w$ Backward wave speed, $w=u_{1}$ $a_{w}$ Gradient of the backward wave cut
$n_{\tilde{c},i},m_{\tilde{c},i}$ Critical accumulation and completion between cut $i$ and cut $i+1$ according to MoC
$n_{\check{c}},m_{\check{c}}$ Critical accumulation and completion of the virtual interception between cut $y$ and cut $z$
$\alpha_{3},\alpha_{2},\alpha_{1}$ Generic coefficients for polynomial FD/MFD, e.g., $G(k)=\alpha_{3}k^{3}+\alpha_{2}k^{2}+\alpha_{1}k^{1}$
Other notations
$L$ Average trip length
$l$ Average lane length
$D$ Total lane length of a network
$r$ Supply disruption magnitude coefficient
$t$ Time
$t_{c}$ Time to reach critical density/accumulation
$t_{\tilde{c},i}$ Time to reach critical density/accumulation between cut $i$ and cut $i+1$
Several assumptions need to be established to define the scope of our work. As the study of fragility involves the recovery process from disruptions, a critical condition to avoid here is the network succumbing to a complete gridlock, where recovery is not possible anymore.